OpenStudy (anonymous):

4x- (6-x)=12+4x solve for x

OpenStudy (anonymous):

4x - 6 + x = 12 + 4x 5x - 6 = 12 + 4x x = 18

OpenStudy (anonymous):

ty, thats what I got!!!

OpenStudy (anonymous):

\[4x-(6-x)=12+4x\\-6+x=12\\x=18\]